the shoe store has twice as many black shoes as it does brown shoes. the total number of shoes is 66. how many brown shoes are there.
Answer by Earlsdon(6103)

Let BL = the number of black shoes and BR = the number of brown shoes.
From the problem statement:
1) BL = 2*BR "the shoe store has twice as many black shoes as it does brown shoes."
2) BL+BR = 66 "the total number of shoes is 66."
Rewrite equation 2) as:
BL = 66-BR and substitute this into equation 1).
(66-BR) = 2*BR Add BR to both sides of the equation.
66 = 3*BR Divide both sides by 3.
22 = BR
There are 22 brown shoes.